# Shear-Rate Dependent Surface Tension of Glass-Forming Fluids

### v1

@@ Abstract
The surface of glass-forming fluids is home to many intriguing dynamical phenomena. In an analogy to surface melting, the physical mechanism conjectured to facilitate ice skating at low temperatures, glasses are covered by layers of enhanced molecular mobility [1, 2]. The highly mobile layers near the surface facilitate the fabrication by layerwise deposition of ultrastable glasses with unique mechanical properties [3, 4, 5], and are a major factor in determining the properties of polymer films [6]. The surface induces a dynamical penetration depth into the bulk that changes non-monotonically with temperature and allows to disentangle the change in mechanisms of relaxation close to a dynamical cross-over temperature $T_{c}$ [7].
@@
We address these questions by molecular dynamics (MD) simulations of a prototypical model of a glass-forming fluid (involving no polymeric or suspension effects), in a simple setup involving a planar surface. Our simulations reveal how an apparent surface tension arises in non-Newtonian fluids that is a mixture of genuine surface effects, and a bulk effect.
Figure 1: Snapshots of the simulation setup without (left) and with (right) imposed shear. Only the central part of the box along $z$ is shown for clarity. We impose ’in-plane’ shear, as shown by the arrows in the coordinate system. Particles are colored according to $\Delta p=p_{z}-(p_{x}+p_{y})/2$ normalized to the interval $[-1,1]$, see Fig. 3.
@@
After equilibration, the simulation box is enlarged to $L_{z}=160\sigma$ keeping the fluid in the center of the box such that around $z=0$, bulk properties are recovered [7]. The interfaces were then relaxed for $250\tau_{0}$, before measurements were performed. Shear flow with rate $\dot{\gamma}$ is imposed in the $(x,y)$-plane tangential to the surface (see Fig. 1) using the SLLOD equations; this is also motivated by the tangential character of the surface tension [19, 20, 21, 22]. Surface tension values were averaged over at least $2000\tau_{0}$. Inspection of the $z$-dependent density profiles revealed no particle-species segregation.
@@
[…]
where the integral extends from the bulk of one phase (the vapor) to the other (the fluid). A factor of $1/2$ is included due to the second interface. Note that convergence of this integral assumes the integrand to vanish at sufficiently large $|z|$, which is the case in the isotropic Newtonian fluid. Essentially we argue that this integral has to be taken with care in the case of a non-Newtonian fluid with shear.
To cross-check our results, we make use of the relation between surface tension and the width of the density profile $\Delta$ [24, 25, 26]:
@@
Figure 3: Pressure differences across the interface: Top panels show $\Delta p_{\alpha\beta}=p_{\alpha}-p_{\beta}$ (blue: $\Delta p_{xy}$; green: $\Delta p_{xz}$; orange: $\Delta p_{yz}$) for temperature $T=0.6$ in (a) the quiescent system and (b) with shear rate $\dot{\gamma}=0.06$. Bottom panels show the relevant pressure difference for the surface tension, $-(\Delta p_{xz}+\Delta p_{yz})/2$, for (c) temperature $T=1$, and (d) $T=0.6$. Color shadings indicate the $z$-intervals identified as bulk and surface regions.
We have determined the surface viscosity by constraining the Green-Kubo integral to particles starting in $z$-layers close to the surface. The notion of “close” can be made precise by looking at the normal-stress differences, $\Delta p_{\alpha\beta}=p_{\alpha}-p_{\beta}$, where $\alpha\beta\in\{x,y,z\}$. Recall that in the isotropic bulk, all $\Delta p_{\alpha\beta}=0$. Close to the surface, the values for $\beta=z$, $\alpha\in\{x,y\}$, deviate from zero since the surface induces an anisotropy. This allows to clearly distinguish a surface layer (top panels in Fig. 3). This surface layer somewhat broadens upon decreasing temperature, consistent with the existence of a growing length scale near the glass transition [7]. We note in passing that the thus identified surface layer to define a surface viscosity (crosses in Fig. 2), indeed follows the empirical relation to the surface tension more closely, although approaching to $T_{c}$ it still appears to break down.
@@
However, the bulk values give non-trivial contributions to the surface-tension integral that are specific to non-Newtonian fluids, and need to be taken into account when determining the surface tension of the sheared fluid. Also this effect becomes more pronounced when temperature is lowered: At $T=1$, the fluid is nearly Newtonian, identified by vanishing normal-stress differences in the bulk (Fig. 3c); here we observe that also the dependence of the pressure anisotropy in the surface region on the shear rate is negligible. But at $T=0.6$, the non-Newtonian effects are already pronounced: with increasing shear rates, increasing bulk normal-stress differences are seen; also the surface regions begin to show a shear rate dependence (Fig. 3d). In Eq. (1), hence an increasing bulk contribution is present the leads to an apparent surface tension in the non-Newtonian fluid.
@@
In summary, we have shown that the surface tension of a typical glass-forming viscoelastic fluid shows a strong shear-rate dependence. In non-Newtonian fluids, normal-stress differences in the bulk and surface anisotropy of the stress tensor conspire to an effective surface tension that is measured using standard techniques that are based on the pressure drop across a liquid film and mix two different contributions. One contribution is a genuine surface contribution, and it comes from the pressure anisotropy in the layers near the surface. In the simulation they can be clearly identified through the different sign of $\Delta p_{\alpha z}$ with respect to the bulk. In addition, there is a non-trivial bulk contribution arising from non-vanishing normal-stress differences, a typical non-Newtonian fluid effect.
@@
This effect might explain why the determination of surface tension values from pressure balances of bulk viscoelastic samples is difficult, as in experiment it will be difficult to disentangle the pure surface from a bulk contribution. Also, hysteresis effects as previsouly reported might be arising from a non-Newtonian bulk contribution: we have monitored the transient evolution of the pressure differences after switching on the shear flow, and found no sign of slow relaxation in the surface contribution. But the bulk quantities are known to exhibit patterns of slow relaxation. Also, a changing volume-to-surface ratio, as in techniques where a fluid droplet is expanded or compresses, might see effects from different mixing of surface and bulk contributions.
@@
Our findings should be relevant to a growing number of techniques measuring interfacial rheological properties in films of complex fluids and even for the printability of bio-inks in surface-tension assisted 3d-printing techniques [35]. For example, biofilms are known to exhibit complex rheology, and characterization of their mechanical surface properties is an important aspect in their growth and removal [36, 37]. The complex interactions in such films can render the surface tension anisotropic [9], and our method of imposing in-plane shear flow in different directions might be a straight-forward way to interrogate the characteristics of the interface in such cases.
@@
On a more theoretical note, it has been shown that the surface of glass-forming fluids reveals non-monotonic changes in the dynamical correlation length governing glassy dynamics [7, 1]. However, this analysis was only based on local density profiles and $z$-resolved density correlation functions. In principle, the discussion of pressure differences as done in our work should give a more natural handle on disentangling surface-induced from bulk effects in such systems. Also, we leave for further discussion the fate of the surface tension in the deeply supercooled regime (below the MCT transition).

### v2

@@ Abstract
The surface of glass-forming fluids is home to many intriguing dynamical phenomena. Glasses are covered by layers of enhanced molecular mobility [1, 2], providing an amorphous analog to surface melting, the physical mechanism conjectured to facilitate ice skating at low temperatures. The highly mobile layers near the surface enable the fabrication by layerwise deposition of ultrastable glasses with unique mechanical properties [3, 4, 5], and are a major factor in determining the properties of polymer films [6]. The surface induces a dynamical penetration depth into the bulk that changes non-monotonically with temperature and allows to disentangle the change in mechanisms of relaxation close to a dynamical cross-over temperature $T_{c}$ [7].
@@
We address these questions by molecular dynamics (MD) simulations of a prototypical model of a glass-forming fluid (involving no polymeric or suspension effects), in a simple setup involving a planar surface. Our simulations reveal how an apparent surface tension arises in non-Newtonian fluids that is a mixture of bulk rheology and genuine surface effects.
Figure 1: Snapshots of the simulation setup without (left) and with (right) imposed shear. Only the central part of the box along $z$ is shown for clarity. We impose ’in-plane’ shear, as shown by the arrows in the coordinate system. Particles are colored according to $\Delta p=p_{z}-(p_{x}+p_{y})/2$ normalized to the interval [1,1][-1,1], see Fig. 3.
@@
After equilibration, the simulation box was enlarged to $L_{z}=160\sigma$ keeping the fluid in the center of the box where around $z=0$, bulk properties are recovered [7]. The interfaces were then relaxed for $250\tau_{0}$, before measurements were performed. Shear flow with rate $\dot{\gamma}$ is imposed in the $(x,y)$-plane tangential to the surface (see Fig. 1) using the SLLOD equations; this is also motivated by the tangential character of the surface tension [19, 20, 21, 22, 23]. Note that our shear protocol keeps the total surface area constant, which is conceptually important for viscoelastic fluids [23]. Surface tension values were averaged over at least $2000\tau_{0}$. Inspection of the $z$-dependent density profiles revealed no particle-species segregation.
@@
$\hat{\sigma}=\frac{1}{2}\int dz\,\left[p_{z}(z)-\frac{1}{2}\left(p_{x}(z)+p_{y}(z)\right)\right]\,$, (1)
where the integral extends from the bulk of one phase (the vapor) to the other (the fluid). A factor of $1/2$ is included due to the second interface in our simulation setup. Note that this integral assumes the integrand to vanish in the bulk regions, which is the case in the isotropic Newtonian fluid. Essentially we argue that this integral has to be taken with care in the case of a non-Newtonian fluid with shear.
To emphasize this point, we also determine the surface tension via the width of the density profile $\Delta$ [25, 26, 27]:
@@
Figure 3: Pressure differences across the interface: Top panels show $\Delta p_{\alpha\beta}=p_{\alpha}-p_{\beta}$ (blue: $\Delta p_{xy}$; green: $\Delta p_{xz}$; orange: $\Delta p_{yz}$) for temperature $T=0.6$ in (a) the quiescent system and (b) with shear rate $\dot{\gamma}=0.06$. Bottom panels show the relevant pressure difference for the surface tension, $(\Delta p_{xz}+\Delta p_{yz})/2$, for (c) temperature $T=1$, and (d) $T=0.6$. Color shadings indicate the $z$-intervals identified as bulk and surface regions.
We have defined a surface viscosity by constraining the Green-Kubo integral to particles starting in $z$-layers close to the surface. The notion of “close” can be made precise by looking at the normal-stress differences, $\Delta p_{\alpha\beta}=p_{\alpha}-p_{\beta}$, where $\alpha,\beta\in\{x,y,z\}$. Recall that in the isotropic bulk, all $\Delta p_{\alpha\beta}=0$. Close to the surface, the values for $\beta=z$, $\alpha\in\{x,y\}$, deviate from zero since the surface induces an anisotropy. This allows to clearly distinguish a surface layer (cf. Fig. 3a), and hence unambiguously define a surface viscosity. This surface layer somewhat broadens upon decreasing temperature, consistent with the existence of a growing length scale near the glass transition [7]. We note in passing that the surface viscosity thus defined indeed follows the empirical relation to the surface tension more closely (crosses in Fig. 2), although approaching to $T_{c}$ a decoupling still appears to occur. We leave this question for a future study.
@@
However, the bulk values give non-trivial contributions to the surface-tension integral that are specific to non-Newtonian fluids, and need to be taken into account when determining the surface tension of the sheared fluid. Also this effect becomes more pronounced when temperature is lowered: At $T=1$, the fluid is nearly Newtonian, identified by vanishing normal-stress differences in the bulk (Fig. 3c); here we observe that also the dependence of the pressure anisotropy in the surface region on the shear rate is negligible. But at $T=0.6$, the non-Newtonian effects are already pronounced: with increasing shear rates, increasing bulk normal-stress differences are seen; also the surface regions begin to show a shear rate dependence (Fig. 3d). In Eq. (1), hence an increasing bulk contribution is present that leads to an apparent surface tension in the non-Newtonian fluid.
@@
In summary, we have shown that the surface tension of a typical glass-forming viscoelastic fluid shows a strong shear-rate dependence. In non-Newtonian fluids, normal-stress differences in the bulk and surface anisotropy of the stress tensor contribute to an effective surface tension that is measured using standard techniques that are based on the pressure drop across a liquid film and mix two different contributions. One contribution is a genuine surface contribution, and it comes from the pressure anisotropy in the layers near the surface. In the simulation they can be clearly identified through the different sign of $\Delta p_{\alpha z}$ with respect to the bulk. In addition, there is a non-trivial bulk contribution arising from non-vanishing normal-stress differences, a typical non-Newtonian fluid effect.
@@
This effect might explain why the determination of surface tension values from pressure balances of bulk viscoelastic samples is difficult, as in experiment it will be difficult to disentangle the pure surface from a bulk contribution. Also, hysteresis effects as previously reported might be arising from a non-Newtonian bulk contribution: we have monitored the transient evolution of the pressure differences after switching on the shear flow, and found no sign of slow relaxation in the surface contribution. But the bulk quantities are known to exhibit patterns of slow relaxation. Also, a changing volume-to-surface ratio, as in techniques where a fluid droplet is expanded or compresses, might see effects from different mixing of surface and bulk contributions.
@@
Our findings should be relevant to various techniques measuring interfacial rheological properties in films of complex fluids [36, 37] and even for the printability of bio-inks in surface-tension assisted 3d-printing techniques [38]. For example, biofilms are known to exhibit complex rheology, and characterization of their mechanical surface properties is an important aspect in their growth and removal [39, 40]. The complex interactions in such films can render the surface tension anisotropic [9], and our method of imposing in-plane shear flow in different directions might be a straight-forward way to interrogate the characteristics of the interface in such cases.
@@
On a more theoretical note, it has been shown that the surface of glass-forming fluids reveals non-monotonic changes in the dynamical correlation length governing glassy dynamics [7, 1]. However, this analysis was only based on local density profiles and $z$-resolved density correlation functions. In principle, the discussion of pressure differences as done in our work should give a more natural handle on disentangling surface-induced from bulk effects in such systems. Also, we leave for further discussion the fate of the surface tension in the deeply supercooled regime (below the MCT transition): eventually, the visco-elastic nature of the glass-forming fluid will be felt, and as one approaches the soft solid state, surface tension and surface energies need to be distinguished [41]. The study of the flow-rate dependence in the glassy regime can probe the different time scales involved in the transition.